### v3

# CodeGRAG: Bridging the Gap between Natural Language and Programming Language via Graphical Retrieval Augmented Generation

Abstract
Utilizing large language models to generate codes has shown promising meaning in software development revolution. Despite the intelligence shown by the general large language models, their specificity in code generation can still be improved due to the syntactic gap and mismatched vocabulary existing among natural language and different programming languages. In this paper, we propose CodeGRAG, a Graphical Retrieval Augmented Code Generation framework to enhance the performance of LLMs. CodeGRAG builds the graphical view of code blocks based on the control flow and data flow of them to fill the gap between programming languages and natural language, which can facilitate natural language based LLMs for better understanding of code syntax and serve as a bridge among different programming languages. To take the extracted structural knowledge into the foundation models, we propose 1) a hard meta-graph prompt template to transform the challenging graphical representation into informative knowledge for tuning-free models and 2) a soft prompting technique that injects the domain knowledge of programming languages into the model parameters via finetuning the models with the help of a pretrained GNN expert model. CodeGRAG significantly improves the code generation ability of LLMs and can even offer performance gain for cross-lingual code generation. Implementation is available at https://anonymous.4open.science/r/Code-5970/ .
CodeGRAG: Bridging the Gap between Natural Language and Programming Language via Graphical Retrieval Augmented Generation
Kounianhua Du1, Jizheng Chen1, Renting Rui1, Huacan Chai1, Lingyue Fu1, Wei Xia2, Yasheng Wang2, Ruiming Tang2, Yong Yu1, Weinan Zhang1 1Shanghai Jiao Tong University, 2 Huawei Noah’s Ark Lab Shanghai, China {kounianhuadu, wnzhang}@sjtu.edu.cn
1 Introduction
In recent years, large language models (LLMs) (Achiam et al., 2023; Touvron et al., 2023a) have shown great impact in various domains. Automated code generation emerges as a captivating frontier (Zheng et al., 2023; Roziere et al., 2023; Shen et al., 2023), promising to revolutionize software development by enabling machines to write and optimize code with minimal human intervention.
However, syntactic gap and mismatched vocabulary among natural language (NL) and programming languages (PL) exist, hindering LLM’s performance on code generation. As illustrated in Figure 1, programming language (marked in blue) contains special tokens such as “int” or “++” that natural language (marked in yellow) doesn’t possess, leading to vocabulary mismatch. Besides, the relations between tokens in programming languages are often structural, e.g., the complex branching and jumps, whereas natural language is arranged simply in sequential manner, leading to syntactic gap. For example, in the control flow graph of the raw code (marked in pink), two “if” blocks (marked in purple) are adjacent and are executed sequentially under certain condition, but they appear to be intervaled in raw textual code.
As discussed above, the innate structures of programming languages are different from that of the sequential-based natural language. The challenges of enhancing a general-purposed large language models for code-related tasks can be summarized into two folds.
Figure 1: Illustration of the gap between the programming language and the natural language.
(C1) How to solve the gap between different languages and better interpret the inherent logic of code blocks. Code, unlike natural language, possesses a well-defined structure that governs its syntax and semantics. This structure provides valuable information about the relationships between different parts of the code, the flow of execution, and the overall organization of the functions Jiang et al. (2021); Guo et al. (2020). General-purpose LLMs regard a code block as a sequence of tokens. By ignoring the inherent structure of codes, they miss out on essential cues that could help them better understand and generate code. In addition, the multi-lingual code generation abilities of LLMs is challenging due to the gap among different programming languages.
(C2) How to inject the innate knowledge of programming languages into general purpose large language models for enhancement. Despite the well representation of the programming knowledge, the ways to inject the knowledge into the NL-based foundation models is also challenging. The structural representation of codes could be hard to understand, which poses a challenge to the capability of the foundation models.
To solve the above challenges, we propose CodeGRAG, a graphical retrieval augmented generation framework for code generation. For (C1), we propose to interpret the code blocks using the composed graph based on the data-flow and control-flow of the code block, which extracts both the semantic level and the logical level information of the code. The composed graphical view could 1) better capture the innate structural knowledge of codes for NL-based language models to understand and 2) model the innate function of code blocks that bridging different programming languages. For (C2), we propose a meta-graph prompting technique for tuning-free models and a soft-prompting technique for tuned models. The meta-graph prompt summarizes the overall information of the extracted graphical view and transforms the challenging and noisy graphical representation into informative knowledge. The soft-prompting technique deals with the graphical view of codes with a pretrained GNN expert network and inject the processed knowledge embedding into the parameters of the general-purpose foundation models with the help of supervised finetuning.
The main contributions of the paper can be summarized as follows:
  • Novel GraphRAG framework for code generation. We propose CodeGRAG that bridges the gap among natural language and programming languages, transfers knowledge among different programming languages, and enhances the ability of LLMs for code generation. CodeGRAG requires only one calling of LLMs and can offer multi-lingual enhancement.
  • Effective graphical view to inform and stimulate the structural programming knowledge of LLMs. We propose an effective graphical view to purify the semantic and logic knowledge from the code space, which offers more useful information than the raw code block and can summarize the cross-lingual knowledge.
  • Effective soft prompting technique to preserve the programming domain knowledge and inject it into LLMs parameters. We propose an effective soft prompting technique, which injects the domain knowledge of programming languages into the model parameters via finetuning LLMs with the assistance of a pretrained GNN expert model.
2 Methodology
Figure 2: Overview of CodeGRAG. (Top) Knowledge Preparation. We extract the control flow and data flow of each external code block and compose them using the read-write signal to obtain the semantic and logical expression of each code block, which is then abstracted into graphical view as hard knowledge document and embedded into GraphEmb as soft knowledge document. The GraphEmb is encoded by a pretrained GNN expert model constrained by the alignment and structure preserving objectives. (Bottem) Retrieval Augmented Generation. We extract query from the task input and retrieve from the external corpus. For tuning free models, we use the hard graphical view to stimulate the structural programming knowledge of LLMs for enhanced generation. For tunable models, we use the soft GraphEmb and inject the programming domain knowledge into LLMs parameters via finetuning them with the GNN expert signals. The expert signals informed LLMs can then produce enhanced generation.
2.1 Overview
In this paper, we leverage both generative models and retrieval models to produce results that are both coherent and informed by the expert graphical knowledge of programming language. The overall process of CodeGRAG is illustrated in Figure 2, which mainly consists of three stages: graphical knowledge base preparation, knowledge querying, and graphical knowledge augmented generation.
2.2 Graphical Knowledge Base Preparation
In this section, we discuss how to extract informative graphical views for code blocks. We analyze the syntax and control information of code blocks and extract their graphical views to better represent the codes. This process can be formulated as, $\forall c_{i}\in D_{\text{pool}}$:
$g_{i}\xleftarrow{}\text{GraphExtractor}(c_{i}),$ (1)
[…]
where $c_{i}$ is the raw code block and $g_{i}$ is the corresponding extracted graphical view.
To capture both the semantic and the logical information, we propose to combine the data flow graph (Aho et al., 2006) and the control flow graph (Allen, 1970) with the read-write signals (Long et al., 2022) to represent the code blocks, both of them are constructed on the base of the abstract syntax tree.
Abstract Syntax Tree (AST). An abstract syntax tree (AST) is a tree data structure that represents the abstract syntactic structure of source code. An AST is constructed by a parser, which reads the source code and creates a tree of nodes. Each node in the tree represents a syntactic construct in the source code, such as a statement, an expression, or a declaration. ASTs have good compactness and can represent the structure of the source code in a clear and concise way.
Data Flow Graph (DFG). The data flow graph (DFG) is a graphical representation of the flow of data dependencies within a program. It is a directed graph that models how data is transformed and propagated through different parts of a program. In DFG, nodes are operands and edges indicate data flows. Two types of edges are considered: 1) operation edges that connect the nodes to be operated and the nodes that receive the operation results; 2) function edges that indicate data flows for function calls and returns. These edges connect nodes, including non-temporary operands and temporary operands, which refer to variables and constants that explicitly exist in the source code, and variables existing only in execution, respectively.
Figure 3: Illustration of the extracted composed syntax graph from the code block. The arrows in the bottom part indicate the names of different edges, which are extracted based on the ASTs.
Control Flow Graph (CFG). The control flow graph (CFG) is a graphical representation of the flow of control or the sequence of execution within a program. It is a directed graph that models the control relationships between different parts of a program. Based on compiler principles, we slightly adjust the design of CFG to better capture the key information of the program. Nodes in CFG are operations in the source code, including standard operations, function calls and returns. Edges indicate the execution order of operations.
Composed Syntax Graph. A composed syntax graph composes the data flow graph and the control flow graph with the read-write flow existing in the code blocks. An illustration of the extracted composed syntax graph is displayed in Figure 3. Different edge types along with their concrete names are given in colors. As for the node names, the middle figure displays the concrete types of nodes (operands) and the right figure displays the properties of nodes.
An illustration of the composed graphical view is in Figure 3. After obtaining the composed syntax graphs, we use them to inform the general-purpose LLMs to bridge the gap between NL and PLs, where both the semantic level and the logic level information are preserved.
2.3 Knowledge Querying
Given a target problem to be completed, we generate informative query of it and use it to retrieve graphical knowledge from the constructed knowledge base.
We extract the problem description of each task to reduce the ambiguity and then concatenate it with the function declaration to serve as the query content, where the functionality and input format of the expected code block are contained. The query of the retrieval includes problem description $Q_{p}$ and function description $Q_{c}$, while each content of the retrieval pool includes raw code block $V_{c}$ and its graphical view $V_{g}$.
To expressively represent the components, we use the encoder $\phi(\cdot)$ of the pretrained NL2Code model to represent the problem description and code snippets. The retrieval function is:
$\mathbf{h^{V}}$ $=\phi(V_{c}\|V_{g}),$ (3)
$\mathbf{h^{Q}}$ $=\phi(Q_{p}\|Q_{c}),$ (4)
Distance $=1-\frac{\mathbf{\mathbf{h^{Q}}\cdot\mathbf{h^{V}}}}{\|\mathbf{h^{Q}}\|\cdot\|\mathbf{h^{V}}\|}.$ (5)
2.4 Graphical Knowledge Augmented Generation
After we obtain the returned graphical view, we inject it to the foundation LLMs for graphical knowledge augmented generation. Since the graphical view is hard to understand, we propose 1) a meta-graph template to transform the graphical view into informative knowledge for tuning-free model and 2) a soft prompting technique to tune the foundation models for their better understanding of the graphical views with the assistance of an expert GNN model.
2.4.1 Hard Meta-Graph Prompt
The original graphical view of a code block could contain hundreds of nodes and edges. A full description of it could cost a overly long context, along with the understanding challenge posed by the long edge lists. Therefore, we propose to use a meta-graph template to abstract the information of the graphical view. The abstracted meta-graph consists of the canonical edge types and node types, which describes the basic topology of the graphical view (Sun and Han, 2013), with the textual features obtained from the ASTs contained in the node and edge features.
Then we use the meta-graph template to transform the retrieved graphical view into digestable knowledge and insert it into the final prompt for generation. As illustrated in Figure 5 in the Appendix, the final prompt consists of three components: the system prompt illustrated in the blue part, the retrieved knowledge and hints illustrated in the green part, and the problem (including task description, function declaration, etc.) illustrated in the yellow part. The three parts are concatenated to be fed into LLMs for knowledge augmented generation.
2.4.2 Soft Prompting with the Expert
Directly hard prompt to the LLMs poses a challenge to the digesting capability of the backbone LLMs, which could fail under the case where the backbone LLMs cannot well understand the graph components. To compress the graphical knowledge into model parameters and help the backbone LLMs to better understand the programming language, we propose a soft prompting technique. The overall procedure can summarized into expert encoding of graphical views, finetuning with the expert signal, and inference.
Expert Encoding of Graphical Views. We design a graph neural network to preserve the semantic and logical information of code blocks. The representation of each node $\mathbf{n}_{i}^{(0)}$ and edge $\mathbf{e}_{i}^{(0)}$ are first initialized with vectors corresponding to the node text and edge text encoded by $\phi_{1}$. A message passing process is first conducted to fuse the semantic and structural information into each node representation.
$\mathbf{m}_{ij}^{(l)}=\mathbf{W}^{(l)}$ $(\mathbf{n}_{i}^{(l-1)}\|\mathbf{e}_{ij}^{(l-1)}),$ (6)
$Q_{j}^{(l)}=$ $\mathbf{W_{Q}}^{(l)}\mathbf{n}_{j}^{(l-1)},$ (7)
$\mathbf{K}_{ij}^{(l)}=\mathbf{W_{K}}^{(l)}m_{ij}^{(l)},$  $\text{\quad}\mathbf{V}_{ij}^{(l)}=\mathbf{W_{V}}^{(l)}\mathbf{m}_{ij}^{(l)},$ (8)
$a_{ij}^{(l)}=\text{softmax}$ (${}_{i\in N(j)}\mathbf{(Q}_{j}^{(l)}\mathbf{K}_{ij}^{(l)}),$ (9)
$\mathbf{n}_{j}^{(l)}=\sum_{i\in N(j)}$ $a_{ij}^{(l)}\mathbf{V}_{ij}^{(l)}.$ (10)
A global attention-based readout is then applied to obtain the graph representation:
$\mathbf{g}=\sum_{i}\text{softmax}(f_{\text{gate}}(\mathbf{n}_{i}^{L}))f_{\text{feat}}(\mathbf{n}_{i}^{L}).$ (11)
The expert encoding network is optimized via the contrastive learning based self-supervised training, which includes the intra-modality contrastive learning and inter-modality contrastive learning. The intra-modality constrastive learning serves for preserving the modality information, while the inter-modality contrastive learning serves for modality alignment.
  • Alignment Contrastive Learning. There are two types of alignment to be ensured: 1) NL-Code (NC) alignment and 2) Code-Graph (CG) alignment. We define the positive pairs for NC alignment purpose as $\mathcal{I}^{+}_{NC}=\{\langle\mathbf{h}_{i}^{V},\mathbf{h}_{i}^{Q}\rangle|i\in D_{\text{train}}\}$ and define the negative pairs for NC alignment purpose as $\mathcal{I}^{-}_{NC}=\{\langle\mathbf{h}_{i}^{V},\mathbf{h}_{j}^{Q}\rangle|i\neq j,i\in D_{\text{train}},j\in D_{\text{train}}\}.$
    And we define the positive pairs for CG alignment purpose as $\mathcal{I}^{+}_{CG}=\{\langle\phi_{1}(c_{i}),\phi_{2}(g_{i})\rangle|i\in D_{\text{train}}\}$ and define the negative pairs for CG alignment purpose as $\mathcal{I}^{-}_{CG}=\{\langle\phi_{1}(c_{i}),\phi_{2}(g_{j})\rangle|i\neq j,i\in D_{\text{train}},j\in D_{\text{train}}\}.$
  • Structure Preserving Contrastive Learning. To preserve the structural information of the graphical views, we perform intra-modality contrastive learning among the graphical views and their corrupted views. Concretely, we corrupt each of the graphical view $g_{i}$ with the edge dropping operation to obtain its corrupted view $g_{i}^{\prime}$. The positive pairs for structure-preserving purpose are then designed as $\mathcal{I}^{+}_{\text{preserve}}=\{\langle\phi_{2}(g_{i}),\phi_{2}(g_{i}^{\prime})\rangle|i\in D_{\text{train}}\}$. The negative pairs for structure preserving purpose are designed as $\mathcal{I}^{-}_{\text{preserve}}=\{\langle\phi_{2}(g_{i}),\phi_{2}(g_{j}^{\prime})\rangle|i\neq j,i\in D_{\text{train}},j\in D_{\text{train}}\}$.
Finetuning with the Expert Soft Signal. To help the backbone LLMs to digest the graphical views, we tune the LLMs with the expert soft signal using supervised finetuning. The prompt for finetuning consists of the system prompt, retrieved knowledge where the expert encoded graphical view is contained using a token embedding, and task prompt, which is illustrated in Figure 5 in the Appendix.
Inference. After the finetuning stage, we used the tuned models to generate codes using the soft prompting template as illustrated in Figure 5 in the Appendix.
3 Experiments
  • RQ1
    Does the proposed CodeGRAG offer performance gain against the base model?
  • RQ2
    Does the proposed graph view abstract more informative knowledge compared with the raw code block?
  • RQ3
    Can soft prompting enhance the capability of the backbone LLMs? Does finetuning with the soft prompting outperforms the simple supervised finetuning?
  • RQ4
    Are the proposed pretraining objectives for the GNN expert effective?
  • RQ5
    What is the impact of each of the components of the graphical view?
  • RQ6
    How is the compatibility of the graphical view?
Table 1: Results of Hard Meta-Graph Prompt on Humaneval-X. (Pass@1)
Model Retrieved Knowledge C++ Python
GPT-3.5-Turbo N/A 57.93 71.95
Code Block (Nashid et al., 2023; Lu et al., 2022) 60.37 72.56
Meta-Graph 62.20 72.56
(Multi-Lingual) Code-Block (Nashid et al., 2023; Lu et al., 2022) 62.20 70.12
(Multi-Lingual) Meta-Graph 64.02 77.44
GPT-4omini N/A 63.41 78.66
Code Block (Nashid et al., 2023; Lu et al., 2022) 65.24 78.66
Meta-Graph 65.85 79.88
(Multi-Lingual) Code-Block (Nashid et al., 2023; Lu et al., 2022) 65.85 79.27
(Multi-Lingual) Meta-Graph 67.07 80.49
Table 2: Results of Soft Prompting. (Pass@1)
Model Finetune CodeForce (C++) APPS (Python)
Gemma 7b N/A 12.83 5.09
SFT 14.76 21.09
Soft Prompting 19.13 26.15
Llama2 13b N/A 9.61 7.29
SFT 11.88 12.06
Soft Prompting 13.62 12.74
CodeLlama 7b N/A 5.20 24.41
SFT 9.87 26.15
Soft Prompting 11.09 30.26
3.1 Setup
In this paper, we evaluate CodeGRAG with the widely used HumanEval-X (Zheng et al., 2023) dataset, which is a multi-lingual code benchmark and CodeForce dataset in which we collect real-world programming problems from codeforces [^1] https://codeforces.com/ website. For CodeForce dataset we include problems categorized by different difficulty levels corresponding to the website and select $469$ problems of difficulty level A for testing. We use greedy decoding strategy to do the generation. The evaluation metric is Pass@1. More details of the retrieval pool and the finetuning setting can be found in Section A in the Appendix.
3.2 Main Results
The main results are summarized in Table 1 and Table 2. From the results, we can draw the following conclusions.
RQ1. The proposed CodeGRAG could offer performance gain against the base model, which validates the effectiveness of the proposed graphical retrieval augmented generation for code generation framework.
RQ2. The model informed by the meta-graph (CodeGRAG) could beat model informed by the raw code block. From the results, we can see that the proposed graph view could summarize the useful structural syntax information and filter out the noises, which could offer more informative knowledge hints than the raw code block. In addition, inserting the intermediate representations of codes into the prompt can stimulate the corresponding programming knowledge of LLMs.
RQ3. From Table 2, we can see that finetuning with the expert soft prompting could offer more performance gain than that brought by simple supervised finetuning. This validates the effectiveness of the designed pretraining expert network and the technique of finetuning with soft prompting, which injects the programming domain knowledge into the LLMs parameters and inform the models with the structural information for gap filling.
3.3 Impacts of the pretraining objectives for the expert GNN (RQ4)
Table 3: Ablation studies on the GNN pretraining losses.
Model Finetune CodeForce (C++) APPS (Python)
Gemma 7b Soft Prompting 19.13 26.15
w/o Alignment 7.88 28.58
w/o Structure-Preserving 11.70 21.50
Llama2 13b Soft Prompting 13.62 12.74
w/o Alignment 11.79 10.76
w/o Structure-Preserving 5.50 11.09
CodeLlama 7b Soft Prompting 11.09 30.26
w/o Alignment 10.92 29.45
w/o Structure-Preserving 10.66 26.59
To study the effectiveness of the proposed pretraining objectives for the expert GNN, we remove each objective to yield different expert GNNs. The results are in Table 3.
From the results, we could see that both the Alignment and the Structure Preserving contribute to the expressiveness of the expert GNN model. The alignment pretraining objective helps to promote the alignment among natural language, programming language, and their graphical views. The structure preserving objective helps to preserve the innate data-flows and control-flows information of code blocks. The two objectives collaborate with each other to yield expressive programming domain knowledge GNN expert model, which encodes external programming knowledge and injects the knowledge into LLMs parameters.
3.4 Impacts of the Components of the Graphical View (RQ5)
In this section, we adjust the inputs of the graphical components to the LLMs. Concretely, we study the information contained in node names, edge names, and the topological structure. The results are presented in Table 4.
Table 4: The impacts of the graph components.
Datasets Python C++
Edge Type Only 73.78 61.59
Edge Type + Node Name 75.00 59.76
Edge Type + Node Type 75.61 59.15
Edge Type + Topological 77.44 64.02
The edge type refers to the type of flows between operands (child, read, write, etc.), the node type refers to the type of operands (DeclStmt, temp, etc.), the node name refers to the name of the intermediate variables, and the topological information refers to the statistics of the concrete numbers of different types of edges. From the results, we can observe that 1) the edge features matter the most in constructing the structural view of code blocks for enhancement, 2) the type of nodes expresses the most in representing operands information, and 3) the overall structure of the graphical view also gives additional information.
3.5 Compatibility Discussion of the Graphical Views(RQ6)
Despite the effectiveness of the proposed graphical views to represent the code blocks, the flexibility and convenience of applying the graphical views extraction process is important for wider application of the proposed method. In this section, we discuss the compatibility of CodeGRAG.
First of all, the extraction process of all the graphical views are front-end. Therefore, this extraction process applies to a wide range of code, even error code. One could also use convenient tools to reformulate the code and improve the pass rate of the extraction process.
In addition, we give the ratio of generated results that can pass the graphical views extraction process, which is denoted by Extraction Rate. The Pass@1 and the Extraction Rate of the generated results passing the graphical extraction process are given in Table 5.
Table 5: The extraction rate of the generated results passing the graphical extraction process.
Generated Codes Pass@1 Extraction Rate
(C++) Code-RAG 62.20 92.07
(C++) CodeGRAG 64.02 92.68
(Python) Code-RAG 71.95 91.46
(Python) CodeGRAG 77.44 96.95
From the results, we could see that the extraction rates are high for codes to pass the graphical views extraction process, even under the situation where the Pass@1 ratios of the generated results are low. This indicates that the application range of the proposed method is wide. In addition, as the code RAG also offers performance gains, one could use multiple views as the retrieval knowledge.
4 Related Work
LLMs for NL2Code. The evolution of the Natural Language to Code translation (NL2Code) task has been significantly influenced by the development of large language models (LLMs). Initially, general LLMs like GPT-J (Radford et al., 2023), GPT-NeoX (Black et al., 2022), and LLaMA (Touvron et al., 2023a), despite not being specifically tailored for code generation, showed notable NL2Code capabilities due to their training on datasets containing extensive code data like the Pile (Gao et al., 2020) and ROOTS (Laurençon et al., 2022). To further enhance these capabilities, additional pretraining specifically focused on code has been employed. PaLM-Coder, an adaptation of the PaLM model (Chowdhery et al., 2023), underwent further training on an extra 7.8 billion code tokens, significantly improving its performance in code-related tasks. Similarly, Code LLaMA (Roziere et al., 2023) represents an advancement of LLaMA2 (Touvron et al., 2023b), benefiting from extended training on over 500 billion code tokens, leading to marked improvements over previous models in both code generation and understanding. These developments underscore the potential of adapting generalist LLMs to specific domains like NL2Code through targeted training, leading to more effective and efficient code translation solutions.
Code Search. The code search methods can be summarized into three folds. Early methods utilizes sparse search to match the query and codes (Hill et al., 2011; Yang and Huang, 2017), which suffers from mismatched vocabulary due to the gap between natural language and codes. Neural methods (Cambronero et al., 2019; Gu et al., 2021) then focus on mapping the query and codes into a joint representation space for more accurate retrieval. With the success of pretrained language models, many methods propose to use pretraining tasks to improve the code understanding abilities and align different language spaces. For example, CodeBERT (Feng et al., 2020) is pretrained on NL-PL pairs of 6 programming languages with the masked language modeling and replaced token detection task. CodeT5 (Wang et al., 2021) supports both code-related understanding and generation tasks through bimodal dual generation. UniXcoder (Guo et al., 2022) integrates the aforementioned pretraining tasks, which is a unified cross-modal pre-trained model. As retrieval augmented generation (RAG) shows its significance in promoting the quality of LLMs generation, works in code RAG start to accumulate. Nashid et al. (2023); Lu et al. (2022) utilize the code blocks as the retrieved knowledge to inform the LLMs with similar code blocks for enhancement. Zhou et al. (2022) uses the programming related document to serve as the retrieval content, injecting auxiliary external programming knowledge into the LLMs generation.
Code Representation. Early methods regard code snippets as sequences of tokens, assuming the adjacent tokens will have strong correlations. This line of methods (Harer et al., 2018; Ben-Nun et al., 2018; Feng et al., 2020; Ciniselli et al., 2021) take programming languages as the same with the natural language, using language models to encode the code snippets too. However, this ignoring of the inherent structure of codes leads to a loss of expressiveness. Methods that take the structural information of codes into consideration then emerge. Mou et al. (2016) used convolution networks over the abstract syntax tree (AST) extracted from codes. Alon et al. (2019) encoded paths sampled from the AST to represent codes. Further exploration into the graphical representation of codes (Allamanis et al., 2017) is conducted to better encode the structures of codes, where more intermediate states of the codes are considered.
5 Conclusion
Despite the expanding role of LLMs in code generation, there are inherent challenges pertaining to their understanding of code syntax. General large language models trained mainly on sequential-based natural language cannot well understand the structural-based programming language, e.g., the branching and jumping in codes. This paper proposes an effective way to build a graphical view of codes to better inform LLMs for code generation. To take the challenging structural graphical knowledge into LLMs, a meta-graph prompt is proposed for tuning-free models and a soft-prompting technique is proposed to inject the structural programming domain knowledge into the parameters of LLMs. By integrating external structural knowledge, CodeGRAG enhances LLMs’ comprehension of code syntax and empowers them to generate code with improved accuracy and fluency.
Limitations
In this paper, we propose a graphical retrieval augmented generation method that can offer enhanced code generation. Despite the efficiency and effectiveness, there are also limitations within this work. For example, dependency on the quality of the external knowledge base could be a potential concern. The quality of the external knowledge base could be improved with regular expression extraction on the noisy texts and codes.
Acknowledgments
This document has been adapted by Emily Allaway from the instructions for earlier ACL and NAACL proceedings, including those for NAACL 2024 by Steven Bethard, Ryan Cotterell and Rui Yan, ACL 2019 by Douwe Kiela and Ivan Vulić, NAACL 2019 by Stephanie Lukin and Alla Roskovskaya, ACL 2018 by Shay Cohen, Kevin Gimpel, and Wei Lu, NAACL 2018 by Margaret Mitchell and Stephanie Lukin, Bib suggestions for (NA)ACL 2017/2018 from Jason Eisner, ACL 2017 by Dan Gildea and Min-Yen Kan, NAACL 2017 by Margaret Mitchell, ACL 2012 by Maggie Li and Michael White, ACL 2010 by Jing-Shin Chang and Philipp Koehn, ACL 2008 by Johanna D. Moore, Simone Teufel, James Allan, and Sadaoki Furui, ACL 2005 by Hwee Tou Ng and Kemal Oflazer, ACL 2002 by Eugene Charniak and Dekang Lin, and earlier ACL and EACL formats written by several people, including John Chen, Henry S. Thompson and Donald Walker. Additional elements were taken from the formatting instructions of the International Joint Conference on Artificial Intelligence and the Conference on Computer Vision and Pattern Recognition.
Appendix A Implementation Details
For the size of retrieval pool, we use 11,913 C++ code snippets and 2,359 python code snippets. Due to the limited access, we do not use a large retrieval corpus for our experiment, which can be enlarged by other people for better performance. We also attach the graph extraction codes for both languages and all other expeirment codes here: https://anonymous.4open.science/r/Code-5970/
For the fintuning details, the learning rate and weight decay for the expert GNN training is 0.001 and 1e-5, repectively. We apply 8-bit quantization and use LoRA for parameter-efficient fine-tuning. The rank of the low-rank matrices in LoRA is uniformly set to 8, alpha set to 16, and dropout is set to 0.05. The LoRA modules are uniformly applied to the Q and V parameter matrices of the attention modules in each layer of the LLM. All the three models are optimized using the AdamW optimizer. For the CodeContest dataset, totally 10609 datapoints are used, and for APPS dataset, 8691 data samples are used to train the model.
Appendix B Prompt Template
Figure 4: Hard meta-graph prompt.
Figure 5: Soft prompting.